Week 50


Machine coordinators Oliver Bruning and Jorg Wenninger





No beam












































Short summary

Link to report from 8:30 meeting

Monday 7th December

Since 00:30: Beam off.
a) Loss of cryo maintain in sector 81 led to dump of stable beams.
b) Electrical perturbation at 04:25 during recovery. Loss of warm compressor in IR4. Magnet trips. Other effects on machine elements.
c) Recovery at around 24:00.
d) Stabel beams as of 2am 4x4 pilots of 6-7e9 but with beam in wrong buckets.

8:30 Summary pptx for December 7

Tuesday 8th December

End of fill around 07:30 after analyzing the issues related to wrong bucket numbers.

Starting one hour fill with LHCb spectrometer OFF.  4 x 4e9 per beam.

Fill aborted due to QPS trigger in Sector 12. Abandon LHCb run and reschedule it for evening or night.

Miscelleanous activities done in parallel:
   1. RF phase adjustment for collision point. (Done by 11am)
   2. Transverde damper noise checks. (Done at )
   3. Test of synchrotron light monitor.

Pre cycle finished and LHCb spectrometer dipole on again. Problems with RF phasing.

Setup for higher intensity: 1 pilots, later n times 2e10

   1. Beam position interlock commissioning.
   2. Dump with beam offset to limit on position interlock (bunched & unbunched).
   3. Injection protection setup.
   4. Abort gap keeper setup (--> requires an access to 2,8 and 6 later in the week).
   5. Review injection sequences.
   6. Collimation check (Q to 1/3 resonance).
   7. Collimation check off-m (fRF trim).
   8. Machine Interlock Test with beam (Ghislain)

Test with synchrotron light monitor?

Ramp tests: 2 x 4e9 per beam.

Starting one hour fill with LHCb spectrometer OFF.  4 x 4e9 per beam.
At end of fill testing of transverese IP steering in IP5.
LHCb run NOT delivered due to nQPS problems in S12. Night spend in trying to establish again circulating beams after end of last ramp.

8:30 Summary pptx for December 8

Wednesday 9th December

Finally managed to establish circulating beams (problem traced to wrong injection settings in triplet orbit correctors in IR2).
Starting one hour fill with LHCb spectrometer OFF.  4 x 4e9 per beam.
At end of fill testing of transverese IP steering in IP5.
Run ended at 11:00 due to lifetime problems in Beam2. CMS could not yet benefit from IP scan but knobs worked fine from machine point of view
LHCb dipole on again.

Preparing Access:
-Point 2, 8 and 6 for LBDS
-Point 4 for RF and BI (undulator)
-Point 5 for TOTEM
-Point 3 for exchange of PRS module in electronics rack for collimator

Continue beam setup for higher intensities. Other miscilleneous work:
-Ramp up of undulator to 400 A in 50 A steps
-Tune feedback for Beam1

ATLAS magnets ramp up (solenoid & toroid)

Stable beam collisions at 450 GeV with 4 x 4 4e9.

8:30 Summary ppt for December 9

LMC Summary pptx for December 9

Thursday 10th December

Test IP transverse adjustment (CMS) and optics verification.
First commissioning of Synchrotron light monitor (mirror adjustments) -> light for Beam2 now available!

Continue beam setup for higher intensities (TCSG/TCDQ Beam1; Transfer line collimators
and TCDIs Beam2; abort gap keeper both beams) -> working on TCDI's Beam1.

Done in parallel:
    -Analysis for losses at MQXA in IR8 during overinjection -> TDI
    -Tune feedback commissioned for Beam1!
    -Injection oscillation correction for Beam1
    -RF phase loop studies
    -Dump test with new ('abort gap keeper' compatible) filling scheme
    -Lifetime optimization for Beam2 by swapping horizontal and vertical tunes
      (avoiding 8kHz excitation near vertical tune)

Suffered a lot from late injection permit from ATLAS after post mortem trigger (ca. 10min to 20min) -> loss of more than an hour in setup time.

Injection test with higher bunch intensity (single bunch).
Optics measurements - verification.

Collimator setup verification with higher intensities (e.g. 2 10^10 / bunch).
Still to come: unsynchronous dump with unbunched beam.

Stable beam collisions at 450 GeV with high bunch intensities: 4 x 2 10^10 per beam<>

8:30 Summary pptx for December 10

17:00 CMS presentation on Vernier Scans ppt

Friday 11th December

9:00:  Transverse IP adjustments in LHCb

10:30: Lost beam (trigger was ATLAS det according to PM on line)
           and struggled to get injection permit after ALICE specrtrometer was powered up

11:20: problem due to the interlocked BPM giving a interlock after the recalibration of BPM

12:30: MKI.B1 generator fault and Beam 1 erratic dump  -> inject Beam2 and perform Q' measurements

14:30: Both beams back -> continue chromaticity measurements in Beam1; swaping tunes for Beam1
           Lifetime increases from 1h to 5h but difficult to measure horizontal tune: many lines in spectrum!

15:00: Trip in MQTL11.L3B2 -> same as yesterday night; caused by faulty current reading (DCCT)
                                                      Faulty signal disconnected -> 16:00 finished pre-cycle for MQTL

16:30: Beams back in machine and start RF studies (1 hour)

19:00: Start transverse IP adjustments in ATLAS, ALICE, CMS (LHCb done in the morning)

21:30: Done with IP adjustements. Use end of fill to measure and adjust chromaticities:
           Old settings from IP adjustment:
           Beam1: Qx = .302; Qy = .272; Qx' = 25.4; Qy' = 7.6; coupling = 0.002; Lifetime 3 h
           Beam2: Qx = .291; Qy = .274; Qx' = 11; Qy' = 10.9; coupling = 0.004; Lifetime   1 h
           New settings after correction:
           Beam1: delta Qx' = -14; delta Qy' = 0
           Beam2: delta Qx' = -5; delta Qy' = -5
           Beam1: Qx = .302; Qy = .272; Qx' = 11;  Qy' = 6.6; coupling = 0.002; Lifetime 25 h
           Beam2: Qx = .299; Qy = .273; Qx' = 4.3; Qy' = 5.5; coupling = 0.004; Lifetime   20 h

00:00: preparing new luminosity fill but lost cryo in sector 81 due to falty temperature reading for RSS

01:30: Beam back

06:30: Lost beam due to PC and QPS system

09:00: Preparing new fill

Rest of the day:
Continue stable beam collisions at 450 GeV with high bunch intensities: 4 x 2 10^10 per beam

Request for acess / time without beam:
-Bernd Dehning to exchange a card for the BLMs (ca. 1 hour)
-ALICE needs one hour of access for fixing TPC crate (ca. 1 hour - compatible with pre-cycle).

8:30 Summary pptx for December 11

Saturday 12th December

Continue stable beam collisions at 450 GeV with high bunch intensities: 4 x 2 10^10 per beam
-Fill at 9am -> 4 hour fill with Beam2 lifetime = 5 hours (Beam1 lifetime = 25 hours)
-Fill at 2pm -> 5 hour fill with Beam2 lifetime = 5 hours (Beam1 lifetime = 25 hours)
-Fill at 8pm -> 2 hour fill with Beam2 lifetime = 12 hours (Beam1 lifetime = 15 hours) Lost fill due to cryo

Studies in parallel:
-Transverse IP adjustments in CMS, LHCb and ALICE
-wire scanners during 'stable beam' mode
-tune adjustments for Beam2

9:00 Summary pptx for December 12

Sunday 13th December

7:00: Recovering from Cryo -> problems related to QPS
                                              -> requires access

18:30: Finished pre-cycle and setting up for beam injection (ca. 20 hours after compressor fault yesterday night).

19:00: Starting Q and Q' adjustements

20:30: Ready for injection but lost cryo due to faulty temperature reading. Need to re-cycle again -> 1.5 hours

22:10: Ready for injection and working on tune and chromaticity adjustments

22:40: start ramp

22:45: lost Beam2 at 800 GeV triggered by BPM interlock but perform measurements on Beam1 at 1.18 TeV

23:30: dump beam and prepare second ramp

02:40: start second ramp: both beams reach 1.18 TeV with goo lifetimes.

04:00: declaring quiet beams for 2 hours

06:00: Dump both beams

Rest of the night: Stable beam with 4 bunches of 2 10^10 per beam.

9:00 Summary pptx for December 13

Monday 14nd December

 Planned machine studies for the next days:

-High intensity at 450 GeV (16 * 5 10^10 per beam)
            separation bumps during injection
            RF adjustments

-Ramp test - pilot collision

-abort gap cleaning

-squeeze and beta* knob tests at 450 GeV (2 hours)



-Orbit bumps at synchrotron light monitor (end of fill)

-intermediate energy dump tests

-TOTEM Roman pot adjustments (end of last Physics fill)

-Detailed check of injection losses on TCDIs and TDI/MQXA, 4 h (anytime late Sunday onwards).

-Injection matching studies, 2-4 h. Lower priority, but if some time spare would like to take some data on injection matching and emittance blowup, to at least start using the techniques we'll have to have working next year.

-Finish up Polarity checks

-Finish up Kick response measurements

-orbit feedback commissioning

-nQPS quench threshold test with beam (< 2e9)

8:30 Summary pptx for December 14